# Program evaluation with remotely sensed outcomes

### v2

@@ G.4 Comparing representations
Figure 14: We contrast optimal versus simple representations of the RSV. The optimal representation combines three predictions. The simple representation uses only one prediction.
@@
Figure 14 visualizes how our optimal representation $\widehat{H}(R)$ compares to the implicit representation of common practice $\textsc{pred}_{Y}(R)$, using the first outcome variable. Even though the satellite image is high dimensional, its representations are scalars. Therefore, we can visualize each village as a point in a plot with $\widehat{H}(R)$ on the vertical axis and $\textsc{pred}_{Y}(R)$ on the horizontal axis. For the sake of visualization, we also plot the binscatter and the best fitting curve. We make such a plot for the second exercise (Figure 14(a); “synthetic samples”) and for the third exercise (Figure 14(b); “real samples”).

### v3

@@ H.4 Comparing Alternative Representations
Figure H.4: We contrast efficient versus simple representations of the RSV. The efficient representation combines three predictions. The simple representation uses only one prediction.
@@
Figure H.4 visualizes how our efficient representation $\widehat{H}(R)$ compares to the implicit representation of common practice $\textsc{pred}_{Y}(R)$, using the first outcome variable. Even though the RSV is high-dimensional, these representations are scalars. Therefore, we can visualize each village as a point in a plot with $\widehat{H}(R)$ on the vertical axis and $\textsc{pred}_{Y}(R)$ on the horizontal axis. For the sake of visualization, we also plot the binscatter and the best fitting curve. We make such a plot for the second exercise (Figure 4(a); “synthetic samples”) and for the third exercise (Figure 4(b); “real samples”).